OpenStudy (anonymous):

Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

y − 4 =3/4 (x + 8) y-4=3/4x+3/4*8 y-4=3/4x+6 -3/4x+y=6+4 −3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y =3/4x + 10 @ivandelgado hope you get it